OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = (-3/4)x-2?

OpenStudy (jdoe0001):

so, what would you say is the slope of -> y = (-3/4)x-2?

OpenStudy (anonymous):

answer choices are: y=(4/3)x-12 y=(3/4)x-5 y=(-4/3)x+20 y=(-3/4)x+13

OpenStudy (anonymous):

that slope would be down three, over four (-3/4), yes?

OpenStudy (jdoe0001):

right, the slope for this -> y = (-3/4)x-2 <-- will be the -3/4

OpenStudy (jdoe0001):

so, perpendicular lines have NEGATIVE RECIPROCAL slopes what does that mean? well, if this line's slope is -3/4 a perpendicular line to it will be RECIPROCAL of -3/4 -4/3 NEGATIVE of that 4/3

OpenStudy (jdoe0001):

so, now you know that the "other line" has a slope of 4/3, and it passes through (12, 4) now use those 2 values in the point-slope form to get your line equation \(\boldsymbol{\huge y-y_1=m(x-x_1)}\) m = slope use your (12, 4) = \(\bf (x_1, y_1)\)

OpenStudy (anonymous):

dude... genius so it would be y=(4/3)x-12, right?

OpenStudy (primeralph):